OpenStudy (anonymous):

Dealing with Calculus rates of change: If a ball is thrown vertically upward with a velocity of 80ft./second, then its height after t seconds is s=80t-16t^2. a) what is the maximum height reached by the ball? b) what is the velocity of the ball when it is 96ft. abov the ground on its way up? On its way down?

OpenStudy (anonymous):

a) Let s' = 0 --> t => max s

OpenStudy (anonymous):

so i set s' equal to zero, then solve for t, correct?

OpenStudy (anonymous):

Yes, s' = 0 --> t!

OpenStudy (anonymous):

ok, I got 5ft./2 seconds. For the second part ,would I substitute 96 in for s'?

OpenStudy (anonymous):

So max height is: s ( 2.5 ) = ....

OpenStudy (anonymous):

ok. Why do you have to set s' to zero?

OpenStudy (anonymous):

I'm really bad with math. Trying to understand this. Ok, so the answer to a) is 2.5 ft./s. How do I tackle question b)?

OpenStudy (anonymous):

a) Let s' = 0 --> t => max s You haven't finished , t = 2.5 sec. Now plug in S ( 2.5) to find max height!

OpenStudy (anonymous):

ok, did that and got 100 ft./s

OpenStudy (anonymous):

40ft/s

OpenStudy (anonymous):

Ok, i took the deriv. of s=80t-16t^2 to get s'=80-32t. set s'=0 to get 0=80-32t, which gave me t=2.5. I plugged my value of t back into my original equation to get 100ft./s : s(2.5)= 80(2.5)-16(2.5)^2= 100ft./s I also plugged it into the s' equation: 80-32(2.5)= 40ft./s

OpenStudy (anonymous):

This is how you show your calculation, not just plug the numbers in: s(2.5)= 16 ( 5*2.5 - 2.5²) = 16 ( 12.5 - 6.25 ) = 16 * 7.25 = ....?

OpenStudy (anonymous):

you used power rule

OpenStudy (anonymous):

Nothing fancy, just plug the numbers in! In order to simplify the calculation, I factorize 16 out. That's it!

OpenStudy (anonymous):

OH!!!!!

OpenStudy (anonymous):

won't the answer still be 100 ft./s? because 2.5^2 is 6.25, and 12.5-6.25 = 6.25. Then 16*6.25= 100

OpenStudy (anonymous):

Did I miss something?

OpenStudy (anonymous):

Oops, I misread my own handwriting! Yes, 16 * 6.25 = 100 ( Sorry!)

OpenStudy (anonymous):

Its ok dude... or dudette! At least I'm paying attention, right? lol

OpenStudy (anonymous):

So, a) Max h = 100 ft

OpenStudy (anonymous):

YAY!!!

OpenStudy (anonymous):

Pay attention to the units t = 2.5 secs, Max height = 100 fts

OpenStudy (anonymous):

i just realized that my answer didn't make sense adding the ft./s on there. You're right, its looking for height

OpenStudy (anonymous):

that the ball is 96 ft. in the air and looking for the velocity while its up in the air and when its falling back down.

OpenStudy (anonymous):

I mean where do you plug 96 in?

OpenStudy (anonymous):

and velocity is speed with direction, and we've got direction, now we need speed. Solving for s?

OpenStudy (anonymous):

Plug in 96ft where s is?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

YAY!!!

OpenStudy (anonymous):

DO i plug 96 into the original equation or the s' equation?

OpenStudy (anonymous):

Plug into S, not S' ( because S is height, S' is velocity )

OpenStudy (anonymous):

I MUST remember that fact. Ok, let me try this

OpenStudy (anonymous):

damn t variables, throwing me off

OpenStudy (anonymous):

S = 96 --> t => V :)

OpenStudy (anonymous):

I got -139,776

OpenStudy (anonymous):

I did something wrong. s(96)= 80(96)-16(96)^2

OpenStudy (anonymous):

No, Plug 96 into where S is ( not where t is): S = 16 ( 5t - t ²) 96 = 16 ( 5t - t ²)

OpenStudy (anonymous):

Can you find t

OpenStudy (anonymous):

that's where I'm stuck

OpenStudy (anonymous):

96 = 16 ( 5t - t ²) 96/16 = 5t - t ² Can you continue?

OpenStudy (anonymous):

\[-t=\sqrt{6-5t}\]

OpenStudy (anonymous):

I assume that t needs to be positive, AND not underneath the radical

OpenStudy (anonymous):

96/16 = 5t - t ² t ² -5t + 6 = 0

OpenStudy (anonymous):

OH!!!!!

OpenStudy (anonymous):

I'm a big stupid :(

OpenStudy (anonymous):

ok, so now t= (2,3)

OpenStudy (anonymous):

now I plug both of my answers back into the original equation and solve for each respectively?

OpenStudy (anonymous):

NO, plug t into S' => V

OpenStudy (anonymous):

ok, and I got.... 16 ft. and -106 ft.

OpenStudy (anonymous):

ft./s

OpenStudy (anonymous):

Can you show your calculation as I showed you?

OpenStudy (anonymous):

s'= 80-32(2)= 80-64=26 ft./s s'= 80-32(2)= 80-96=-106 ft./s

OpenStudy (anonymous):

whoops, its 1, NOT 26

OpenStudy (anonymous):

oomg, 16

OpenStudy (anonymous):

V (3) = ...?

OpenStudy (anonymous):

V(3)= s'= 80-32(3)= 80- 96=-16

OpenStudy (anonymous):

-16 ft.s

OpenStudy (anonymous):

Yup, now we arrive the same result this time =)

OpenStudy (anonymous):

V(3) = -16 ft/sec

OpenStudy (anonymous):

YAY!!!! danke shon Chlorophyll!!! you are very patient! I GREATLY appreciate your help!!!!

OpenStudy (anonymous):

gracias, thank you!!! Whichever language you like hearing "thank you" in

OpenStudy (anonymous):

well, at least anything in spanish, german, or english

OpenStudy (anonymous):

i took half that time, for one question. I know I'm difficult, but I don't try to be

OpenStudy (anonymous):

I did, thank you very, very much. I've got to go study anatomy. I hope that you'll be around to help again, soon. You have a good night/day, depending where you're located

OpenStudy (anonymous):

Good luck with your quizzes, test, exams, .... :)